find three consecutive even integers such that the square of the largest exceeds the sum of the squares of the other two by 12
Answer by mananth(16949)

x, x+2, x+4
(x+4)^2= x^2+(x+2)^2 +12
x^2+8x+16=x^2+x^2+4x+4+12
x^2-4x=0
x(x-4)=0
x=0 OR x=4
0,2,4
4,6,8
m.ananth@hotmail.ca